UT20090706 LUCIFER Commissioning

Observers: WSeifert, NAgeorges (LBT), DThompson (Tucson)
Telescope Operator: D Gonzalez Huerta
Software Support: CBiddick
Instrument Support:
Telescope Support:

Summary

Details (UT unless otherwise noted)

04:00 opening dome. Clouds have largely cleared off, RH~65%

05:30 Having trouble with the EL encoders not switching back to strip mode after a slew (even manually). TSargent called. Related to IT#2242? Prior to this, LUCIFER was seeing a large (~60 arcsec) offset to the pointing model.

05:50 Having trouble acquiring a 13.7 mag star. DGH checked outside and there are (patchy) clouds.

05:59 TSargent calls back. It looks like a "Dan issue", according to Tom. This is a new failure.

We hit a +Z limit on the primary once, then later hit the +Y limit twice. Later still (10:15) we hit a +X limit and then shortly thereafter we it the +RXRY limit.

On acquisition image #13 GCS claimed to have found a star, but there is nothing there. Are we still falling back to old (pre-SExtractor) algorithms?

12:00 Stop for flats

-- DavidThompson - 05 Jul 2009
I Attachment Action Size Date Who Comment
daq.20090706.gifgif daq.20090706.gif manage 50 K 08 Jul 2009 - 17:38 JohnHill Plot of Telescope and Mirror Temperatures
Topic revision: r6 - 08 Jul 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback